Three oranges cost half of what nine apples cost. If one orange and one apple together cost 30 cents how much does one orange cost.
This question is not from a textbook-the instructor gave it to us.
A+O=.30 THEN A=.30-O
3O=9A/2
3O=9(.30-O)/2
3O=(2.7-9O)/2
3O=(2.7-9O)/2
6O=2.7-9O
60+9O=2.7
15O=2.7
O=2.7/15
O=.18 FOR EACH ORANGE.
A=30-18
A=12.
PROOF
3*.18=9*.12/2
.54=1.08/2
.54=.54

Found 2 solutions by checkley75, lammy60:
Answer by checkley75(3666)

A+O=.30 THEN A=.30-O
3O=9A/2
3O=9(.30-O)/2
3O=(2.7-9O)/2
3O=(2.7-9O)/2
6O=2.7-9O
60+9O=2.7
15O=2.7
O=2.7/15
O=.18 FOR EACH ORANGE.
A=30-18
A=12.
PROOF
3*.18=9*.12/2
.54=1.08/2
.54=.54

Answer by lammy60(2)


3O=9A/2
3O=9(.30-O)/2
3O=(2.7-9O)/2
3O=(2.7-9O)/2
6O=2.7-9O
60+9O=2.7
15O=2.7
O=2.7/15
O=.18 FOR EACH ORANGE.
A=30-18
A=12.
PROOF
3*.18=9*.12/2
.54=1.08/2
.54=.54
Feel free to use the tutors here again and again!